solve the equation.
1/x+1 + 1/x-1 =2/x^2-1

Answer by stanbon(75887)

solve the equation.
1/(x+1) + 1/(x-1) =2/(x^2-1)
Multiply thru by x^2-1 to get:
x-1 + x+1 = 2
2x=2
x=1
---------
But x cannot be 1 as that would make two of the original denomiators zero.
---------
Conclusion: No solution
===============
Cheers,
Stan H.